Evaluate the logarithmic equation for three values of x that are greater than 1, three values of x that are between 0 and 1, and at x=1. Use the resulting ordered pairs to plot the graph. State the equation of the line asymptotic to the graph (if any).
y= log5x

I just wanted to know if the work below was correct? I want to make sure I have it correct before turning it in. Thank you.
y=log5x
x=0.1:
y=log(5*0.1)
y=log(0.5)
highlight(y=-0.3)
x=1, y=0.7
x=2, y=1
x=3, y=1.18
x= 4, y = 2
x= .5, y=-1
x= .25, y= -2
x= .125, y = -3

Answer by stanbon(75887)

y=log5x
x=0.1:
y=log(5*0.1)
y=log(0.5)
highlight(y=-0.3)
x=1, y=0.7
x=2, y=1
x=3, y=1.18
---
You are good up to this point.
Then, the wheels came off.
x = 4, y = log(5*4) = 1.301
x = .5, y = log(5*0.5) = 0.3979
x = .25, y = log(5*0.25) = 0.09691...
x = 0.125, y = log(5*0.125) = -0.2041...
=================
Cheers,
Stan H.
=================

x= 4, y = 2
x= .5, y=-1
x= .25, y= -2
x= .125, y = -3